OpenStudy (anonymous):

Evaluate and simplify.... 63/3*2x(x-4) for x =7

OpenStudy (anonymous):

First, plug in x. \[{63\over 3} \cdot 2(7)((7)-4)\] Can you do the rest?

OpenStudy (anonymous):

I can give it a shot, its the only way I will learn.... Just a quick second...

OpenStudy (anonymous):

147?

OpenStudy (anonymous):

Nope. Show your steps. Did you go from left to right? \[{63\over 3}\ \cdot \ 2 \ \cdot \ (7) \ \cdot \ (7-4)\]

OpenStudy (anonymous):

Yes, I did.. Ok let me look at this

OpenStudy (anonymous):

I am still getting 147

OpenStudy (anonymous):

\[ {63\over 3}\ \cdot \ 2 \ \cdot \ (7) \ \cdot \ (7-4) \\ = 21\ \cdot \ 2 \ \cdot \ (7) \ \cdot \ (7-4) \\ = 42 \ \cdot \ (7) \ \cdot \ (7-4) \\ = 294 \ \cdot \ (7-4) \\ = 294 \ \cdot \ 3 \\ = 882\]

OpenStudy (anonymous):

I see.. I think the only way to work these problems is to learn how to properly work out the problems just like you just did. I thank you very much for helping me.